OpenStudy (anonymous):

A baseball team celebrated their championship by having a party. The team members ate spaghetti and meatballs and drank sodas. If the team ordered 15 bowls of spaghetti that cost $4.95 each, and 15 sodas that cost $1.29 each, how much did the party cost? A. $74.25 B. $54.90 C. $93.60 D. $19.35

OpenStudy (anonymous):

15 * $4.95 + 15 * 1.29 = 74.25 + 19.35 = 93.60

OpenStudy (anonymous):

The party costs $93.60

OpenStudy (anonymous):

thanks